OpenStudy (anonymous):

Elias hits a ball into the air. The equation h=-16t^2+60t+3 models the hight (h) in feet of the ball after t seconds. I don't really understand how long the ball is in the air.

OpenStudy (anonymous):

wat we have to find out...

OpenStudy (anonymous):

how long the ball is in the air

OpenStudy (anonymous):

Can u recheck ur question...i feel something is missing...

OpenStudy (anonymous):

ok, so a girl his a ball into the air... the equation h=-16t^2+60t+3 models the height (h) in feet of the ball after t seconds. How long is the ball in the air?

OpenStudy (anonymous):

i meant to say "hits" a ball.

OpenStudy (anonymous):

srry. i have to go.

hartnn (hartnn):

the ball will remain in the air till the height becomes zero,that is when it hits the ground.....so put h=0 .......then u will get a quadratic equation in t which solved gives u time t....can u solve that quadratic equation?

OpenStudy (anonymous):

yeah, one sec...

OpenStudy (anonymous):

hmm. i think i did something wrong. i know the quadratic equation, but i ended up so far with: -60 plus or minus the square root of 3792 divided by -32.

hartnn (hartnn):

do u know any other method of solving quadratic equation?? because factorising is much simpler here.....

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so, i have to find what minus 60 will give me -48 now, right?

OpenStudy (anonymous):

60-12=48. but 12 - 60= -48!

hartnn (hartnn):

can u use calculator? then u can find root of 3792 easily and what u had done becomes more simpler because of calculator.....

OpenStudy (anonymous):

-16t^2+12t+48t+3, right?

hartnn (hartnn):

nopes....12+48=60 but 12*48 does not equal 3*16.... but can u use calculator?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the square of 3792 is 61.6, rounded

hartnn (hartnn):

ok so in -60 plus or minus the square root of 3792 divided by -32. now u have (-60-61.6)/(-32) ......solve this using calcy...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-.05,3.8!

hartnn (hartnn):

since time cannot be negative u have t=3.8 so the ball will be in the air for 3.8 seconds :)

OpenStudy (anonymous):

right, because 3.8 is the only sensible answer to the problem! that was my last problem! thanks so much, and have a great school year!!

hartnn (hartnn):

:D i completed my school years ago......

OpenStudy (anonymous):

oh :)

OpenStudy (anonymous):

bye

hartnn (hartnn):

bye :)